OpenStudy (♥ashley♥):

A cone has a diameter of 24 in. and a height of 10 in. What is the volume of the cone? Use 3.14 as an approximation for π. Express your final answer in tenths.

Mehek (mehek14):

do you know the formula?

OpenStudy (calculusxy):

The formula for finding the volume of a cone is: V = Bh

OpenStudy (calculusxy):

B = area of the base

OpenStudy (♥ashley♥):

1/3 π × r^2 × h

OpenStudy (calculusxy):

Sorry yes. I confused it with the cylinder

Mehek (mehek14):

or \(πr^2\dfrac{h} {3}\)

Mehek (mehek14):

so what do you have as the radius and height?

OpenStudy (calculusxy):

So it's just: \[\frac{ B \times h }{ 1/3 }\]

OpenStudy (♥ashley♥):

The height is 10. And the diameter 24 divided by 2 gives us the radius of 12.

Mehek (mehek14):

ok so \(π12^2\dfrac{10} {3}\)

OpenStudy (♥ashley♥):

It equals 40

Mehek (mehek14):

the whole thing?

Mehek (mehek14):

\(3.14 * 12^2 = 452.16\)

Mehek (mehek14):

now multiply \(452.16*\dfrac{10}{3}\)

OpenStudy (♥ashley♥):

It equals 1507.2

OpenStudy (calculusxy):

yes

Mehek (mehek14):

yea

OpenStudy (♥ashley♥):

Thanks everyone!

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (calculusxy):

Welcome :)